Chae Corporation uses the weighted-average method in its process costing system. This month, the beginning inventory in the first processing department consisted of 670 units. The costs and percentage completion of these units in beginning inventory were:

Cost Percent Complete
  Materials costs $7,600       70%            
  Conversion costs $6,600       15%            

A total of 9,800 units were started and 8,350 units were transferred to the second processing department during the month. The following costs were incurred in the first processing department during the month:

  Materials costs $137,600
  Conversion costs $323,600

The ending inventory was 85% complete with respect to materials and 70% complete with respect to conversion costs.

Note: Your answers may differ from those offered below due to rounding error. In all cases, select the answer that is the closest to the answer you computed. To reduce rounding error, carry out all computations to at least three decimal places.

The cost per equivalent unit for materials for the month in the first processing department is closest to:
Answer #1

Ending inventory Unit = 670+9800-8350 = 2120

Equivalent unit of material = 8350+(2120*85%) = 10152

Cost per equivalent unit of material = (7600+137600)/10152 = 14.30

The cost per equivalent unit for materials for the month in the first processing department is closest to: $14.30